OpenStudy (anonymous):

The current in a solenoid with 23 turns per centimeter is 0.45A . The solenoid has a radius of 1.6cm . A long, straight wire runs along the axis of the solenoid, carrying a current of 15A . Part A Find the magnitude of the net magnetic field a radial distance of 0.80cm from the straight wire.

OpenStudy (anonymous):

B=uNI/L B=UNI/2r B=UI/2pir U=4pi x10^-7 Fm=qvBsin

OpenStudy (whpalmer4):

What will be the field contributed by the straight wire alone?

OpenStudy (anonymous):

what do you mean?

OpenStudy (whpalmer4):

Well, the current through the wire is going to generate a magnetic field, isn't it? The net magnetic field will be the linear combination of the field due to the solenoid and the field due to the wire.

OpenStudy (anonymous):

yes

OpenStudy (whpalmer4):

So, do you know how to find the field contributed by the straight wire? That should be the easier of the two parts...

OpenStudy (anonymous):

really sad answer, but no.

OpenStudy (whpalmer4):

Damn, now I have to figure it out :-)

OpenStudy (anonymous):

and another equation is Fm=LIBsin

OpenStudy (whpalmer4):

for that I think the formula is \[B = \frac{\mu_0 I}{2\pi r}\]

OpenStudy (anonymous):

I was going to use that, but I thought it would only be is it was circular because of the 2pi r

OpenStudy (anonymous):

i got 1.8 x10^-8

OpenStudy (whpalmer4):

Well, the magnetic field lines make circles around the wire. If you are using your right hand and pointing your thumb along the wire in the direction of the current, the field lines follow your fingers, going in the direction they point.

OpenStudy (anonymous):

oh yea, sorry every time it says straight conductor I think the field lines are going straight, bad mistake, sorry.

OpenStudy (anonymous):

I got i got 1.8 x10^-8

OpenStudy (whpalmer4):

Hmm, I got 0.000375 for my answer

OpenStudy (anonymous):

what did you use for r and I?

OpenStudy (whpalmer4):

\[B = \frac{4\pi * 10^{-7}*15}{2\pi *0.8*0.01} = \]

OpenStudy (whpalmer4):

I = 15 A, r = 0.8 cm * 1 m/100 cm= 0.008 m

OpenStudy (anonymous):

ok, I think your right. I made a mistake in my calculations

OpenStudy (anonymous):

go on

OpenStudy (whpalmer4):

so now we need to find the field inside the solenoid due to the solenoid (r = 1.6 cm > 0.8 cm)

OpenStudy (anonymous):

ok, so B=UI/2pir equation

OpenStudy (whpalmer4):

for that, the formula appears to be \[B = \mu_0(\frac{N}{L})I\]

OpenStudy (anonymous):

or that . I really don't like magnetism.

OpenStudy (whpalmer4):

here I = 0.45 A and N/L = 23/1 cm or 23 / 0.01m

OpenStudy (whpalmer4):

yeah, I'm not crazy about it either :-)

OpenStudy (anonymous):

8.1x10^-8 ??????

OpenStudy (whpalmer4):

but these field lines are running parallel to the straight wire. I'm not really sure how to combine them...

OpenStudy (whpalmer4):

can you show me how you got that number?

OpenStudy (anonymous):

{(4pix10^-7)(23 turns)/ (1.6)(100)}(0.45)

OpenStudy (whpalmer4):

why are you including the radius? the field due to the solenoid is uniform within the core of the solenoid...

OpenStudy (anonymous):

do i put 0.80 cm instead ?

OpenStudy (whpalmer4):

http://www.pa.msu.edu/courses/1997spring/PHY232/lectures/ampereslaw/solenoid.html no, there's no radial component to worry about, just\[ B = \mu_0(\frac{N}{L})I\]

OpenStudy (anonymous):

but what would L be?

OpenStudy (whpalmer4):

N/L is turns per length. 23 turns in 1 cm or 23 turns in 0.01 m

OpenStudy (whpalmer4):

I have to leave now, sorry I can't be of more help!

OpenStudy (anonymous):

ok. Could you tell me what to do after I get the magnetic fields

OpenStudy (anonymous):

I got 1.30 x10^-3